OpenStudy (anonymous):

Simplify: (8x^3)/(2 cubedroot x)

OpenStudy (anonymous):

\[4x ^{8/3}\]

OpenStudy (anonymous):

how'd you get that? :/

OpenStudy (anonymous):

you can use this prperty of exponents: (a^m)/(a^n)=a^m-n

OpenStudy (anonymous):

First divide the coefficients so 8/2 = 4 then x^3 divde by cubedroot x is same as \[x ^{3-3/2}\]

OpenStudy (anonymous):

Making the coefficient 8/3? (I'm no good with fractions)

OpenStudy (anonymous):

sorry if I confused you... :D I meant divide the numbers so 8 over 2 is 4

OpenStudy (anonymous):

ahh, lol. but i just dont get how to get the coefficient to 8/3